OpenStudy (anonymous):

A flower vase, in the form of a hexagonal prism, is to be filled with 512 cubic inches of water. Find the height of the water if the wet portion of the flower vase and its volume are numerically equal.

OpenStudy (anonymous):

V = SA = 512 V = ( 3 √3 /2 ) a² * H 512 = 2.6 a² * H ->H = 197.1 / a² SA = 2.6 a² + 6a h 512 = 2.6 a² + 6a ( 197.1 / a²) = 2.6 a² + 1,182.6 / a 512a = 2.6a³ + 1,182.6 so just put that in your calc. and you will have your answer.

OpenStudy (anonymous):

Hi, been working this pen pen and paper, and with the internet, I have this as the best solution to the above problem: http://ph.answers.yahoo.com/question/index?qid=20100505051316AAXv8po though, this is isn't my work, I have checked this and the result is magnificent.

OpenStudy (anonymous):

V = SA = 512 V = ( 3 √3 /2 ) a² * H 512 = 2.6 a² * H ->H = 197.1 / a² SA = 2.6 a² + 6a h 512 = 2.6 a² + 6a ( 197.1 / a²) = 2.6 a² + 1,182.6 / a 512a = 2.6a³ + 1,182.6 sorry aboutthe first one.

OpenStudy (anonymous):

well im on my school laptop so can you tye it out on here?

OpenStudy (anonymous):

your answer is 34.88 inches.

OpenStudy (anonymous):

the composite area shown consists of a semicircle an a region bounded by line BC and the two congruent arcs AB and AC whose radii are each 12 inches and each central angle is 60˚. Find the area of this figure.